### v1

@@ 6 Beltrami Optimization On Spherical Topology (BOOST)
7 Implementation
For the training detail of the SBN model, pleaase refer to [xu2025neural]. In downstream applications, besides manually designing the weight of each loss term, we may treat this multi-loss optimization as a multitask learning problem. To balance the optimization of multiple loss terms, we adapted the method from [liebel2018auxiliary] with modifications. Specifically, we introduced learnable loss weights $c_{i}$ to dynamically adjust the emphasis on each task. Loss terms below predefined thresholds were excluded from optimization to focus on more challenging tasks, preventing the model from overfitting to simpler objectives. This approach is detailed in Alg.2. All experiments were conducted on an NVIDIA A40 GPU with 44GB memory, ensuring efficient computation.
Algorithm 2 multi-task loss with threshold
1:
@@ 7 Implementation
    Output $f$ and input $\mu$, $p$ of the SBN and parameters $(\theta,\sigma,t)$ in post-composed mapping.
@@
    loss functions $\mathcal{L}_{i}$, corresponding thresholds $\mathcal{N}_{i}$ and learnable loss weights $c_{i},(i=1,2,...,n)$
2:loss $\mathcal{L}_{total}$
[…]
4:for $i=1$ to $n$ do
5:  Compute $\mathcal{L}_{i}=\mathcal{L}_{i}(f,\mu,p,\theta,\sigma,t)$
6:  if $\mathcal{L}_{i}>\mathcal{N}_{i}$ then
7:   $\mathcal{L}_{total}=\mathcal{L}_{total}+c_{i}\mathcal{L}_{i}$
8:  end if
9:end for
8 Experimental results
Numerical experiments were conducted to evaluate the efficacy of our framework BOOST, which composed of several applications in landmark, intensity, and hybrid registration problems, encompassing both realistic and synthetic scenarios.
8.1 Cortical Surface Registration
To further demonstrate the versatility and effectiveness of our framework, we conducted a series of experiments focusing on the spherical parameterization of cortical surfaces. We tested Alg.1 by performing a landmark alignment task on the spherical parameterization of right cortical surface data. We followed the experiment setup from FLASH [choi2015flash]. On each cortical surface, there are six landmark curves, including Central Sulcus (CS), Inferior Frontal Sulcus (IFS), Superior Frontal Sulcus (SFS), Inferior Temporal Sulcus (ITS), Superior Temporal Sulcus (STS) and Postcentral Sulcus (PostCS). We first used FLASH [choi2015flash] to obtain a spherical conformal parameterization of each cortical surface and aligned landmarks on the pair of spherical cortical surfaces. Mathematically, denote six landmark curves on the spherical parameterization of two cortical surfaces by $S_{1,i},S_{2,i}(i=1,\ldots,6)$ and similarity measure functions of landmark pairs to be matched by $d_{sim}$. Our goal is to find $f\colon\mathbb{S}^{2}\rightarrow\mathbb{S}^{2}$ such that
  1. 1.
    minimizing $d_{sim}(S_{1,i},S_{2,i})$ for all $i=1.\ldots,6$;
  2. 2.
    $f$ is diffeomorphic;
@@ 8.1 Cortical Surface Registration
$\mathcal{L}_{lm}(f)=\mathcal{L}_{task}(f)+\mathcal{L}_{reg}(f),$
@@
$\mathcal{L}_{task}(f)=\frac{1}{6}\sum\limits_{i=1}^{6}d_{sim}(f(S_{1,i}),S_{2,i}).$
For the regularization terms involved in included but not limited to this experiment, we usually use a weighted sum of $\mathcal{L}_{bm},\mathcal{L}_{folding},\mathcal{L}_{bs},\mathcal{L}_{\mu}$ and $\mathcal{L}_{sm}$.
In prior work [choi2015flash], $d_{sim}$ is nodewise L2 square distance and this is indispensible for them to linearize the optimization problem via the variational method. To satisfy this, Choi et.al [choi2015flash] used resampling to ensure that the number of nodes in the corresponding curves is the same. We applied the same resampling operation, followed by a rotation to minimize the average landmark L2 distance. The registration was then performed using our SBN-based framework, and the results were compared with those of FLASH[choi2015flash] in terms of both angle distortion and landmark matching error. Numerical results show that our method achieves competitive, or even superior, landmark matching accuracy meanwhile always yielding consistently lower angle distortions. Fig.3 and Fig.4 compare the performance of different methods in an example of registering brain 32 to brain 35, in terms of visual effect and statistics of induced Beltrami coefficients. See Tab.1 for a more comprehensive comparison.
Figure 3: Registration results for brain 32 to brain 35 using six sulcal landmarks. Landmarks are grouped (Group 1: CS, ITS, STS, postCS; Group 2: IFS, SFS) for visual clarity. Columns, from left to right: source/target landmark pairs (green: moving surface, blue: template), followed by results from FLASH[choi2015flash], our method with resampling, and our method with chamfer distance. Red shows deformed sulci, blue shows targets.
@@
Brains 35 and 32 0 0.118 0.0017 0.0015 0 0.103 0.0009 0.0011 0 0.114 0.0031 0.0009
Brains 22 and 28 0 0.407 0.0038 0.0004 0 0.098 0.0030 0.0036 0 0.258 0.0021 0.0002
Brains 29 and 13 1 0.493 0.0031 0.0030 0 0.108 0.0037 0.0043 0 0.119 0.0054 0.0009
Brains 8 and 18 5 0.548 0.0047 0.0083 0 0.085 0.0042 0.0047 0 0.110 0.0079 0.0018
Brains 32 and 2 0 0.372 0.0037 0.0075 0 0.103 0.0030 0.0032 0 0.116 0.0028 0.0006
Table 1: Comparison of landmark matching performance and angle distortion among FLASH[choi2015flash] and BOOST (with two variations: method 1—with landmark resampling; method 2—using chamfer distance as loss, without landmark resampling). For each pair of brains, the table provides the number of overlap failures, average $|\mu|$, mean squared distance, and mean chamfer distance. Lower values are better in all columns except number of overlaps.
Moreover, in both experimental and real-world brain landmark matching tasks, it is impractical to require that corresponding curves from different brains have the same number of points. Curves like the Central Sulcus of different brains may be represented by different numbers of sampled points due to variations in brain size, shape, or imaging resolution. This discrepancy poses a significant challenge for traditional alignment methods that rely on point-to-point comparisons. Additionally, if the same curves in two subjects indeed have very different lengths, as the following large deformation case indicates, these resampling methods may not accurately capture true anatomical correspondence, potentially introducing errors and artifacts, affecting the reliability of subsequent analyzes.
@@
For clarity, we denote the former application of our framework by method 1, and this by method 2. A result is shown in the rightmost column of Fig.3. As Tab. 1 shown, Replacing the L2 nodewise distance by the sum of the chamfer distance and endpoint loss consistently achieves visually better landmark matching performance and smaller angle distortion than FLASH [choi2015flash].
@@
In addition, even when the initial landmark distortion is small, FLASH [choi2015flash] may fail to achieve a diffeomorphic deformation when imposing too many landmark constraints, whereas our method remains robust. For example, when registering brain 8 (moving surface) to brain 18 (template surface), FLASH is able to handle registering ITS and STS, or CS, SFS, IFS, and postCS separately, but cannot provide a diffeomorphic registration when all six sulci are involved. In contrast, our method can simultaneously register all six landmark curves, and the registration accuracy for both ITS and STS, as well as CS, SFS, IFS, and postCS, exceeds that of FLASH even when FLASH handles these sulci in individual smaller groups. As shown in Fig.5, compared to FLASH, we can match the landmark curve well without inducing too much distortion in other regions. For example, the ITS-STS registration result by FLASH is highly biologically implausible, whereas ours induces much less distortion in regions away from landmark curves, no matter whether using landmark resampling or not.
Figure 5: Registration of brain 8 to brain 18 with 6 sulcal landmarks(for visualization and color conventions, see Fig. 3). Note that in this figure, first row is visualization of CS, IFS, SFS and postCS, while the second row is visualization of ITS and STS. The second column are two independent results done by FLASH [choi2015flash], respectively.
@@
Our another advantage is that it remains robust in the large deformation case. See Fig.6 for an example registering CS, IFS and SFS of brain 39 to those of brain 35. As shown, there are big differences between lengths and positions between two curves in each landmark pair. FLASH [choi2015flash] fails to give a non-overlap deformation while ours still provides a diffeomorphic registration result with little distortion.
@@
Figure 6: Registration of brain 39 to brain 35 using CS, IFS, SFS as landmark curves (for color conventions, see Fig.3). The first column is visualization on the original moving face. The second and third ones are results using method 1 and 2, respectively.
@@
Furthermore, our method is seamlessly compatible with geometric feature alignment in anatomical parcellation. For illustration, we registered 16 subjects’ right-hemi cortical surface in the OASIS dataset [lamontagne2019oasis] to population averages from FreeSurfer[fischl2012freesurfer]. More detailedly, initial spherical representations for each subject’s cortical surface were generated using FreeSurfer. We designated the fsaverage6 surface, a standard population-average spherical template distributed with FreeSurfer, as the common registration target space. For the parcellation registration, we used FreeSurfer to generate parcellation labels (i.e. 34 cortical areas) for each subject including the fsaverage6 surface. A preprocessing step involves mapping subject-specific geometric features (e.g., sulcal depth, curvature, or other relevant anatomical descriptors) from the individual’s native spherical parameterization onto the fsaverage6 template. This is achieved by interpolating these feature values at the vertex locations defined by the fsaverage6 mesh. This procedure ensures that feature information from the moving subject is represented on the common template grid, establishing vertex-wise correspondence for comparative analysis and guiding the alignment.
Let $\Phi$ be the non-rigid deformation, and $M,F$ be geometric features on the moving sphere and the fixed sphere,respectively. During optimization, the feature similarity between $M\circ\Phi$ and $F$ is estimated using Pearson correlation coefficient(PCC),
@@
where $P$ denotes the number of parcels, and the parcellation atlas used in registration are all generated by FreeSurfer.
@@
Again, we applied BOOST with $\mathcal{L}_{task}=4\cdot d_{par}+d_{corr}$ for the experiment. The parcellation accuracy was quantified using the Dice score (higher values denote greater overlap). The alignment of geometric features was evaluated by comparing each subject’s sulcal depth to the FreeSurfer fsaverage atlas. Similarity was measured via PCC, with higher PCC indicating better alignment. Our method achieves 0.8844 PCC for sulci depth on average and 0.9680 Dice score on average, and Fig.8.3 shows a more detailed result for 34 regions parceled by the atlas provided by FreeSurfer. Note that to the best of our knowledge, there have not been works conducting the same experiments on the OASIS3 dataset before, and the aim of this experiment is to demonstrate the flexibility and capacity of our method and its potential to be widely applied in many tasks. So, it is not our intention to focus on comparing the performance with others. However, it is worth mentioning that in SUGAR [ren2024sugar], Ren et al. conduct the same experiment on the ADRC dataset, which also consists of brain data from Alzheimer Disease subjects, and their method achieved 0.839 Dice score and 0.904 PCC. This implies that our method is also competitive and potentially superior to other mainstream algorithms.
8.2 Extreme Distortion
We further challenge our method by conducting a synthetic and highly demanding landmark-matching experiment designed to create extreme distortions. Specifically, we randomly selected $2n(n=2,3,4)$ groups of landmarks $(A_{i},B_{i},C_{i},D_{i})_{i=1,2,...,2n}$, distributed evenly on both hemispheres. Each group of landmarks is constrained that adjacent points are mapped to each other, thereby constructing $2n$ large-scale twists on the spherical surface. The first row of Fig.7 displays representative pairs of landmarks, while the task loss in this setup is defined as
$\mathcal{L}_{task}(\mu,p)$ $=\frac{1}{8n}\sum\limits_{i=1}^{n}(|\mathcal{F}(\mu,p)(A_{i})-B_{i}|^{2}+|\mathcal{F}(\mu,p)(B_{i})-C_{i}|^{2}$
$\quad\quad\quad\quad+|\mathcal{F}(\mu,p)(C_{i})-D_{i}|^{2}+|\mathcal{F}(\mu,p)(D_{i})-A_{i}|^{2}).$
Fig.7 is a result of $n=4$ and the average L2 squared error of each pair is decreased to $7.96\times e^{-4}$ after optimization. As demonstrated by the qualitative and quantitative results, SBN is highly adept at dealing with tasks requiring extremely large deformation mappings: bijectivity is rigorously preserved since all Beltrami coefficients retain a norm less than 1 throughout, while smoothness and conformality are maintained by constraining both $|\mu|$ and $|\nabla\mu|$. Thus, even under extreme landmark-driven warping, regions outside the immediate deformation zones remain minimally distorted.
Figure 7: Results for the extreme distortion experiment 8.2. Top row: Landmarks to be matched, forming large-scale twists. Middle row: Deformation result of the surface after matching. Bottom row: Mapping results shown for both hemispheres, with color indicating the magnitude of the Beltrami coefficient.
8.3 Hybrid Matching
Lastly, we tested our method in a synthetic experiment as well as a real-scenario application related to cortical surface registration. In the first test, the method was applied to a synthetic task requiring a diffeomorphic deformation that morphs an ’I’-shaped region on the sphere into a ’C’ shape, while matching six pairs of corresponding landmarks and the entire image intensity. As shown in Fig.8.3, the deformed image closely resembles the target image, landmarks are consistently matched and the obtained registration remains bijective.
@@ 8.3 Hybrid Matching
Next, we tested our algorithm to register mean curvature and six sulci of two cortical surfaces. We used brain 32 as moving surface and brain 35 as template surface, and matched the six pairs of landmark curves meanwhile maximizing the correlation of their mean curvature. Fig.8 is the distribution of mean curvature on cortical surfaces, which shows that distribution of moving surface is more similar to that of fixed surface after deformation. Notably, after hybrid matching, the similarity in mean curvature distribution is significantly improved—from a pre-registration correlation coefficient of 0.083 to 0.581 post-registration (as opposed to just 0.123 when performing only landmark matching). We also compare the landmark matching effect of these two methods in Fig.9, and hybrid matching delivers both appropriate landmark registration and mean curvature alignment, enhancing the accuracy and anatomical interpretability of surface correspondence, which is an advantage in quantitative neuroimaging analysis and medical diagnosis.
[ Sulci Depth and Parcellation Registration Result] [Uncaptioned image]
[Synthetic Experiment of deforming an ’I’ shape region on the sphere to the ’C’.shape] [Uncaptioned image]
@@
Figure 8: Mean curvature correlations between the moving surface and the target surface before and after registration. From left to right: (1) before deformation; (2) after hybrid matching registration; (3) after only-landmark matching registration. Alignment is visually and statistically improved after hybrid matching.
@@
Figure 9: Comparing registration results: (left) fixed cortical surface, (center) deformed moving surface after hybrid matching, (right) deformed moving surface using landmark-only matching. Hybrid matching achieves both precise landmark alignment and improved mean curvature correlation.

### v2

@@ 6 Beltrami Optimization On Spherical Topology (BOOST)
7 Experiments and Applications
Numerical experiments were conducted to evaluate the efficacy of our framework BOOST, which comprises several applications in landmark, intensity, and hybrid registration problems, encompassing both realistic and synthetic scenarios. For the training detail of the SBN model, please refer to [xu2025neural]. In downstream applications, all experiments related to BOOST were conducted on an NVIDIA A40 GPU with 44GB memory, ensuring efficient computation, and the numerical algorithms were implemented using MATLAB on an 8-core Intel i7 machine.
7.1 Experiments
We first tested our methods on two synthetic experiments. In the first test, the method was applied to a hybrid registration task requiring a bijective deformation that morphs an ’I’-shaped region on the sphere into a ’C’ shape, while matching six pairs of corresponding landmarks and the entire image intensity. As shown in Figure 3, the deformed image very closely resembles the target image, landmarks are consistently matched and the obtained registration remains bijective.
We further challenged our method by conducting a landmark-matching experiment involving extreme distortions. Specifically, we randomly selected 2n(n=2,3,4)2n\ (n=2,3,4) groups of landmarks $(A_{i},B_{i},C_{i},D_{i})_{i=1,2,...,2n}$, distributed evenly on both hemispheres. Each group of landmarks is constrained that adjacent points were mapped to each other, thereby constructing $2n$ large-scale twists on the spherical surface. The top-left picture of Figure 4 displays representative pairs of landmarks, while the task loss in this setup is defined as
$\mathcal{L}_{task}(f)=\frac{1}{8n}\sum\limits_{i=1}^{n}(|f(A_{i})-B_{i}|^{2}+|f(B_{i})-C_{i}|^{2}+|f(C_{i})-D_{i}|^{2}+|f(D_{i})-A_{i}|^{2}).$
The Top-right picture of Figure 4 shows the final registration grid for a case that $n=4$ and the average L2 squared error of each pair is decreased to $7.96\times 10^{-4}$ after optimization, and the second row is the Beltrami distribution of the result mapping. As demonstrated by the qualitative and quantitative results, SBN is highly adept at dealing with tasks requiring extremely large deformation mappings: bijectivity is rigorously preserved since all Beltrami coefficients retain a norm less than 1 throughout, while smoothness and conformality are maintained by constraining both $|\mu|$ and $|\nabla\mu|$. Thus, even under extreme landmark-driven warping, regions outside the immediate deformation zones remain minimally distorted.
Figure 3: Deformation from an ’I’-shape region to a ’C’-shape region. The moving surface (showing a vertical yellow strip) is transformed to align with the target surface (displaying a C-shaped yellow region). The third panel shows the successfully registered result after applying the transformation. The rightmost visualization presents the registration grid.
Figure 4: Results for the extreme distortion experiment. Top row: Landmarks to be matched, forming large-scale twists. Middle row: Deformation result of the surface after matching. Bottom row: Mapping results shown for both hemispheres, with color indicating the magnitude of the Beltrami coefficient.
7.2 Applications to brain surface registration
Cortical surface registration plays a crucial role in aligning functional and anatomical features across individuals. Due to its requirement on the high registration accuracy, preserving topology, and minimizing registration distortion, the problem is challenging. To demonstrate the effectiveness of our framework, we conducted a series of experiments focusing on the registration between spherical parameterizations of cortical surfaces.
We tested Algorithm 1 by performing a landmark alignment task on the spherical parameterization of right cortical surface data. We followed the experiment setup from FLASH [choi2015flash]. On each cortical surface, there are six landmark curves, including Central Sulcus (CS), Inferior Frontal Sulcus (IFS), Superior Frontal Sulcus (SFS), Inferior Temporal Sulcus (ITS), Superior Temporal Sulcus (STS) and Postcentral Sulcus (PostCS). We first used FLASH [choi2015flash] to obtain a spherical conformal parameterization of each cortical surface and aligned landmarks on the pair of spherical cortical surfaces. Mathematically, denote six landmark curves on the spherical parameterization of two cortical surfaces by $S_{1,i},S_{2,i}(i=1,\ldots,6)$ and similarity measure functions of landmark pairs to be matched by $d_{sim}$. Our goal is to find $f\colon\mathbb{S}^{2}\rightarrow\mathbb{S}^{2}$ such that
@@ 7.2 Applications to brain surface registration
    minimizing $d_{sim}(f(S_{1,i}),S_{2,i})$ for all $i=1.\ldots,6$;
@@
    $f$ is bijective;
@@
$\mathcal{L}_{\text{lm}}(f)=\mathcal{L}_{\text{task}}(f)+\mathcal{L}_{\text{reg}}(f),$
@@
$\mathcal{L}_{\text{task}}(f)=\frac{1}{6}\sum\limits_{i=1}^{6}d_{\text{sim}}(f(S_{1,i}),S_{2,i}).$
For the regularization terms involved in included but not limited to this experiment, we usually use a weighted sum of $\mathcal{L}_{\text{bm}},\mathcal{L}_{\text{folding}},\mathcal{L}_{\text{bs}},\mathcal{L}_{\text{BC}}$ and $\mathcal{L}_{\text{smooth}}$.
In prior work [choi2015flash], $d_{sim}$ is nodewise L2 square distance and this is indispensable for them to linearize the optimization problem via the variational method. To satisfy this, Choi et al. [choi2015flash] used resampling to ensure that the number of nodes in the corresponding curves is the same. We applied the same resampling operation, followed by a rotation to minimize the average landmark L2 distance. The registration was then performed using our SBN-based framework, and the results were compared with those of FLASH[choi2015flash] in terms of both angle distortion and landmark matching error. The hyperparameter $\lambda$ in FLASH was set to 3 and the weights of $\mathcal{L}_{\text{task}},\mathcal{L}_{\text{bm}},\mathcal{L}_{\text{folding}},\mathcal{L}_{\text{bs}},\mathcal{L}_{\text{BC}}$ and $\mathcal{L}_{\text{smooth}}$ were 5, 1, 20, 0.5, 0.1 and 0.01, respectively. Numerical results show that our method achieves competitive, or even superior, landmark matching accuracy meanwhile always yielding consistently lower angle distortions. Figure 5 and Figure 6 compare the performance of different methods in an example of registering brain 32 to brain 35, in terms of visual effect and statistics of induced Beltrami coefficients. See Table 1 for a more comprehensive comparison.
Figure 5: Registration results for brain 32 to brain 35 using six sulcal landmarks. Landmarks are grouped (Group 1: CS, ITS, STS, postCS; Group 2: IFS, SFS) for visual clarity. Columns, from left to right: source/target landmark pairs (green: moving surface, blue: target surface), results from FLASH[choi2015flash], results of our method with resampling, and results of our method with chamfer distance. Red curves are the deformed sulci, and blue ones shows target sulci.
@@
Brains 35 and 32 0 0.123 0.0018 0.0014 0 0.09122 0.0007 0.0009 0 0.09126 0.0027 0.0007
Brains 29 and 13 44 0.296 0.0034 0.0030 0 0.091 0.0018 0.0012 0 0.098 0.0056 0.0004
Brains 18 and 8 107 0.307 0.0022 0.0029 0 0.096 0.0011 0.0012 0 0.074 0.0117 0.0007
Brains 21 and 39 128 0.366 0.0411 0.0632 0 0.158 0.0018 0.0020 0 0.144 0.0061 0.0009
Brains 10 and 9 157 0.387 0.0048 0.0037 0 0.118 0.0015 0.0014 0 0.113 0.0056 0.0003
Table 1: Comparison of landmark matching performance and angle distortion among FLASH[choi2015flash] and BOOST (with two variations: method 1—with landmark resampling; method 2—using chamfer distance as loss, without landmark resampling). For each pair of brains, the table provides the number of overlap failures, average $|\mu|$, mean squared distance, and mean chamfer distance. Lower values are better in all columns.
Moreover, in both experimental and real-world brain landmark matching tasks, it is impractical to require that corresponding curves from different brains have the same number of points. Curves like the Central Sulcus of different brains may be represented by different numbers of sampled points due to variations in brain size, shape, or imaging resolution. This discrepancy poses a significant challenge for traditional alignment methods that rely on point-to-point comparisons. Additionally, if the same curves in two subjects indeed have very different lengths, as the following large deformation case indicates, these resampling methods may not accurately capture true anatomical correspondence and the traditional methods may fail.
@@
The weights of $\mathcal{L}_{\text{task}},\mathcal{L}_{\text{bm}},\mathcal{L}_{\text{folding}},\mathcal{L}_{\text{bs}},\mathcal{L}_{\text{BC}}$ and $\mathcal{L}_{\text{smooth}}$ in this method were the same as before. For clarity, we denote the former application of our framework by method 1, and this by method 2. A result is shown in the rightmost column of Figure 5. As shown in Table 1, replacing the L2 nodewise distance by the sum of the chamfer distance and endpoint loss consistently achieves visually better landmark matching performance and smaller angle distortion than FLASH [choi2015flash].
@@
In addition, even when the initial landmark distortion is small, FLASH [choi2015flash] may fail to achieve a bijective deformation when imposing too many landmark constraints, whereas our method remains robust. For example, when registering brain 8 (moving surface) to brain 18 (template surface), FLASH is able to handle registering ITS and STS, or CS, SFS, IFS, and postCS separately, but cannot provide a non-overlap registration when all six sulci are involved. In contrast, our method can simultaneously register all six landmark curves, and the registration accuracy for all six sulci, exceeds that of FLASH even when FLASH handles these sulci in individual smaller groups. But for a fair comparison of the capacity of controlling geometric distortion, in Figure 7 we compare the registration result when all methods are required to register only CS, SFS, IFS, and postCS. Compared to FLASH, we can match the landmark curve well without inducing too much distortion in other regions. As shown in Figure 7, the registration result by FLASH is highly biologically implausible, while ours induce much less distortion in regions away from landmark curves, no matter whether using landmark resampling or not.
Figure 7: Registration result comparison of brain 8 to brain 18 with 4 sulcal landmarks in two views (for visualization and color conventions, see Figure 5).
@@
Another advantage of BOOST is that it remains robust in the large deformation case. See Figure 8 for an example registering CS, IFS and SFS of brain 39 to those of brain 35. As shown, there are big differences between lengths and positions between two curves in each landmark pair. FLASH [choi2015flash] fails to give a non-overlap deformation while ours still provides a bijective registration result with little distortion. To demonstrate robustness with complex landmark configurations, we performed registration on 11 sulci pairs annotated on two left hemisphere cortical surfaces, including the lateral sulcus, central sulcus, postcentral sulcus, intraparietal sulcus, transverse occipital sulcus, inferior frontal sulcus, precentral sulcus, superior temporal sulcus, inferior temporal sulcus, superior frontal sulcus, and lateral occipital sulcus. While FLASH failed to maintain bijectivity under this dense landmark constraint, both proposed methods successfully produced non-overlapping mappings with accurate landmark correspondence (see Figure 9).
@@
Figure 8: Registration of brain 39 to brain 35 using CS, IFS, SFS as landmark curves (for color conventions, see Figure 5). The first column is visualization on the original moving face. The second and third ones are results using method 1 and 2, respectively.
Figure 9: Visualization of registration results by our two methods. Left: Original landmark configuration before registration shown on spherical parameterizations (View 1 and View 2), where blue indicates target landmarks and green indicates moving surface landmarks. Method 1 (top row) and Method 2 (bottom row): Registration results displayed on spherical representations and mapped to the actual cortical surface. In both methods’ results, red landmarks represent the deformed moving surface landmarks after registration, while blue landmarks denote the target positions.
@@
Furthermore, our method is seamlessly compatible with alignment of geometric feature and anatomical parcellation. For illustration, we registered 189 subjects’ right-hemi cortical surface in the OASIS3 dataset [lamontagne2019oasis] to population averages from FreeSurfer[fischl2012freesurfer]. More specifically, initial spherical representations for each subject’s cortical surface were generated using FreeSurfer. We used the fsaverage6 surface, a standard population-average spherical template provided by FreeSurfer, as the common registration target space. For the parcellation registration, we used FreeSurfer to generate parcellation labels (i.e. 35 cortical areas) for each subject including the fsaverage6 surface. Let $\Phi$ be the non-rigid deformation, and $M,F$ be geometric features on the moving sphere and the fixed sphere, respectively. During optimization, the similarity of features between $M\circ\Phi$ and $F$ is estimated using the Normalized correlation coefficient(NCC),
@@
where $P$ denotes the number of parcels, and the parcellation atlas used in registration was generated by FreeSurfer.
@@
We applied BOOST with $\mathcal{L}_{task}=10\cdot d_{par}+d_{corr}$ for the experiment. The parcellation accuracy was quantified using the Dice score (higher values denote greater overlap). The alignment of geometric features was evaluated by comparing each subject’s sulcal depth to the FreeSurfer fsaverage atlas. Performance was measured by two metrics, NCC and Dice score, with higher NCC values indicating better alignment accuracy in sulcal depth and higher Dice score meaning greater parcellation similarity. The mean and standard deviation of NCC and Dice score are 0.9122 ± 0.0169 and 0.9216 ± 0.0262, respectively. Figure 10 shows a more detailed result for 35 regions parceled by the atlas provided by FreeSurfer. It is worth mentioning that in SUGAR [ren2024sugar], Ren et al. conduct the same experiment on the ADRC dataset, which also consists of brain data from Alzheimer Disease subjects, and their method achieved 0.904 NCC and 0.839 Dice score. This implies that our method is also competitive and potentially superior to other mainstream algorithms.
@@
Lastly, we applied our algorithm to registering sulcal depth and six sulci of two cortical surfaces. We used the left cortical surface of subject 30001 and 30003 in the OASIS3 dataset as the moving surface and target surface, respectively, and matched the six pairs of landmark curves meanwhile maximizing the correlation of their sulcal depth. Figure 11 is the distribution of sulcal depth on cortical surfaces, which shows that distribution of sulcal depth on the moving surface is more similar to that on the fixed surface after deformation. Notably, after hybrid matching, the similarity in sulcal depth distribution is improved from a pre-registration correlation coefficient of 0.490 to 0.705 post-registration (as opposed to just 0.166 when performing only landmark matching). We also compare the landmark matching effect of these two methods in Figure 12, and hybrid matching delivers both appropriate landmark registration and sulcal depth alignment, enhancing the accuracy and plausibility of surface correspondence, which is an advantage in quantitative neuroimaging analysis and medical diagnosis.
Figure 10: Quantitative Evaluation of Surface Registration Accuracy for Parcellation and Sulcal Depth Data. The left panels show parcellation registration results with Dice scores, indicating regional alignment accuracy across different brain parcels. The right panels present sulci depth registration results with correlation coefficients, measuring the correspondence of sulcal depth patterns between registered surfaces and the averaged template.
@@
Figure 11: Sulcal depth correlations between the moving surface and the target surface before and after registration. From left to right: (1) before deformation; (2) after hybrid matching registration; (3) after only-landmark matching registration. Alignment is visually and statistically improved after hybrid matching.
@@
Figure 12: Comparing registration results with sulcal landmarks of the moving surface highlighted in red and the target landmarks highlighted in blue: (left) fixed cortical surface, (center) deformed moving surface after hybrid matching, (right) deformed moving surface using landmark-only matching. Hybrid matching achieves both precise landmark alignment and improved sulcal depth correlation.